OpenStudy (anonymous):

Help me solve this trig equation. I'll write it down in equation form in the next post. By the way the domain for the function is between 0 and 2pi (closed interval).

OpenStudy (anonymous):

\[-2\cos (x)-2(\cos(x))^2+2(\sin(x))^2\] Need to set this equal to zero and solve for x.

OpenStudy (anonymous):

(sin(x))2 becomes sin^2x and with the Pythagorean identity you have \[-2\cos(x)-2(\cos ^{2}x)+2(1-\cos ^{2}x)\] That's the first step. From there it's just algebra. I'm at work so I can't write the whole solution out, but if you you get stuck, post how far you got and someone can take a look at it.

OpenStudy (anonymous):

ok cool thanks I'm sure I can do it now